# When More is Less: Understanding Chain-of-Thought Length in LLMs

### v1

@@ 7 Conclusion
[…]
Appendix A Additional Real world Experiment Details
A.1 Implementation Details
In real world experiments, we investigate how the optimal CoT length varies between these two models and with different question difficulties. To create solutions with varying step lengths, we follow [Fu et al., 2023] by using in-context examples (8-shots) with three different levels of complexity to guide the model in generating solutions with different step counts. For each set of in-context examples, we sample 20 times, resulting in a total of 60 samples per question.
[…]
When evaluating the results, questions with accuracy $<0.01$ or $>0.99$ (indicating all incorrect or all correct responses) are excluded, as their accuracy does not vary with step length changes.
A.2 Task Difficulty v.s. Optimal CoT Lengths
To further investigate the relationship between task difficulty and optimal CoT lengths on real world datasets, we conduct experiments on different models. The results (Figure 7 and 8) are impressive that results on all models show a significant correlation between the task difficulties and optimal lengths.
@@ A.1 Implementation Details
Appendix B Synthetic Arithmetic Problem Discussions
B.1 Contrast to vanilla arithmetic problem
B.2 Same Subtask Difficulty
A Chain-of-Thought (CoT) containing sub-tasks of varying difficulty can be seen as a mixture of CoTs with the same difficulty level. On the other hand, allowing CoTs to generate sub-tasks of different lengths increases the overall difficulty of CoTs of the same length, making them harder to study. Third, under Assumption E.1, convexity analysis shows that the final accuracy function (Proposition 5) is concave. Therefore, to maximize accuracy, all sub-tasks should have the same difficulty level.
Appendix C Subtask Loss
@@
Appendix D Synthetic Experiment Details
Appendix E General Error Functions
Assumption E.1.
@@ Appendix E General Error Functions
Appendix F Proof
@@ F.1 Proof of Proposition 5
We begin the proof by incorporating the notation from F.2. We have
@@
Let $F(x^{*}(T),T)=g^{\prime}(x^{*}(T))=0,\forall T.$ We want to see how $x^{*}(T)$ changes as $T$ changes, therefore we take total derivative w.r.t. $T$. By the chain rule,
@@
According to Lemma F.1, $F(x_{0}(T),T)<0$. Since $F(x^{*}(T),T)=0,\text{ and }\,F_{x}\,\Bigl{(}x^{*}(T),\,T\Bigr{)}<0,$ we have $x_{0}(T)>x^{*}(T)$.
@@
F.5 Technical Lemmas
Lemma F.1.

### v3

@@ C.1 Implementation Details
More Details of Figure 2(b). When evaluating the results, questions with accuracy $<0.01$ or $>0.99$ (indicating all incorrect or all correct responses) are excluded, as their accuracy does not vary with step length changes.
To better understand the reliability of the observed trend between task difficulty and optimal Chain-of-Thought (CoT) length, we compute a 95% confidence interval around the linear regression line. Specifically, we use standard methods based on the Student’s t-distribution to estimate uncertainty in the predicted values. The confidence band reflects how much the estimated mean CoT length is expected to vary given the finite sample size and the distribution of data points.
C.2 More Experimental Results
Additional results for Figure 2(b). To further investigate the relationship between task difficulty and optimal CoT lengths on real world datasets, we conduct experiments on different models. The results (Figure 6 and 7) are impressive that results on all models show a significant correlation between the task difficulties and optimal lengths.
@@ C.2 More Experimental Results
Results on MMLU STEM dataset. We also conduct experiments on the MMLU STEM dataset using the Qwen2.5 Series instruct models under the same settings as the MATH dataset. The results, shown in Figures 8 and 9, exhibit similar trends to those observed on the MATH dataset.
(a) Optimal CoT length vs. Model size (Qwen2.5 series).
(b) Optimal vs. Longest CoT length accuracy on MMLU STEM dataset.
Figure 8: Real-world CoT length observations. (a) Larger models tend to achieve optimal performance with shorter CoTs. (b) Accuracy for CoTs of optimal length is significantly higher than that of the longest CoTs.
(a) Qwen2.5-1.5B-Instruct
(b) Qwen2.5-7B-Instruct
(c) Qwen2.5-32B-Instruct
(d) Qwen2.5-72B-Instruct
Figure 9: Evaluation between task difficulties and optimal CoT lengths on MMLU STEM datasets with Qwen2.5 Series Instruct models.
Appendix D Supplementary Details on Real World Experiment for RL Simplicity Bias
For Figure 1(b), we use Qwen2.5-7B-Instruct [Qwen et al., 2025] as the base model, Group Relative Policy Optimization with R1-like prompting [Shao et al., 2024, Guo et al., 2025] for the reinforcement learning process, and LeetCode-2K [Xia et al., 2025] as the training and evaluation dataset. We take the following training configuration by default:
Table 1: Hyperparameter settings for real-world RL experiments with Qwen2.5-instruct models.
Learning Rate Max Epochs Rollout Samples Reverse KL Coefficient Entropy Loss Coefficient Effective Batch Size
5e-7 10 16 1e-3 5e-3 256
Appendix E Additional Synthetic Experiment Details
E.1 Training details
E.2 Observation of subtask loss
@@
Appendix F Theoretical Results under Broader Scenarios
F.1 General Errors
In the simple case we discussed in Section 4.1, we discussed the trend of overall accuracy with respect to $N$ and the variation of optimal $N$ with $M$ and $T$, assuming the subtask error rate is a linear function. In the following discussion, we aim to derive conclusions corresponding to more general error rate functions. We find that as long as the error function satisfies some basic assumptions on the monotonicity and convexity of the error functions, the above conclusions still hold.
Assumption F.1.
@@ F.1 General Errors
With Assumption F.1 and F.2), the core insights from the linear case can be generalized.
{restatable}
theoremgeneral For a noise function $10<\sigma(T)<$ 1 and a subtask error rate function $10<E(N,M,T)<$ 1 satisfying Assumptions F.1 and F.2, the general final accuracy function $A(N)$ from Proposition 1 has the following properties:
  • $\lim_{N\rightarrow+\infty}A(N)=0$. (Excessively long chains always fail.)
  • If $A(N)$ has a maximum at $N^{*}>1$, then $N^{*}$ has a lower bound related to $M$ and $T$:
    $N^{*}\geq N_{LB}(M,T)=E_{N}^{-1}\left(1-\frac{1}{e^{2}(1-\sigma(T))};M,T\right),$ (6)
    where $E_{N}^{-1}(\cdot;M,T)$ is the inverse of $E(N,M,T)$ with respect to $N$.
The monotonicity of $E_{N}^{-1}$ with respect to $M$ (decreasing) and $T$ (increasing, assuming $\sigma(T)$ doesn’t dominate adversely) implies that the qualitative scaling laws (Corollaries stemming from Theorem 4.1) still hold under general conditions, supporting the empirically observed Simplicity Bias and the inverted U-shaped performance.
Corollary F.3.
As the model becomes stronger, $E^{-1}$ decreases monotonically with respect to $M$, which leads to a decrease of $N(M,T)$.
Corollary F.4.
As the task becomes harder, $E^{-1}$ is monotonically increasing with respect to $T$, which leads to an increase in $N(M,T)$.
F.2 Random Error
In Theorem 4.1 and F.2, we make a strong assumption that all sub-question or sub-answer errors are identical, which does not align well with real-world scenarios. In practice, each sub-task may exhibit a different error rate. However, they generally follow a trade-off: the more the task is decomposed, the easier each sub-task becomes. Specifically, we can model the error rate of each sub-task as a random variable with a fixed expectation that monotonically decreases with the number of CoT steps $N$.
To simplify the problem, here we assume $\sigma_{i}\sim B(\alpha_{1}(T),\beta_{1}(T))$ to be the sub-question error rate, and $e_{i}\sim B(\alpha_{2}(N,M,T),\beta_{2}(N,M,T))$ to be the sub-answer error rate. Then, as a variant of Proposition 1, the expectation of final accuracy is $\mathbb{E}\left[\prod_{i=1}^{N}(1-e_{i})(1-\sigma_{i})\right]$.
It is worth noting that each $\sigma_{i}$ or $e_{i}$ is not independent. If most steps are easy (i.e., have low error rates), the remaining steps are more likely to be easy as well. Moreover, if a particular step serves as a self-validation step, its high accuracy can influence the correctness of other steps that depend on it. This also provides an interpretation for reasoning models exhibiting backtracking behavior.
{restatable}
theoremrandom Let $\alpha_{1}=T$, $\beta_{1}=C-T$, $\alpha_{2}=T$, and $\beta_{2}=NM-T$. Then the expected error rates for sub-questions and sub-answers are given by $\mathbb{E}[\sigma_{i}]=\frac{T}{C}$ and $\mathbb{E}[e_{i}]=\frac{T}{MN}$, respectively. Based on these estimates, we can derive an upper bound $\hat{A}(N)$ on the final accuracy
$\mathbb{E}\left[\prod_{i=1}^{N}(1-e_{i})(1-\sigma_{i})\right]\leq\hat{A}(N)=\left[\left(1-\frac{T}{C+2N-1}\right)\left(1-\frac{T}{NM+2N-1}\right)\right]^{N},$
which initially increases and then decreases as the number of CoT steps $N$ grows.
This suggests that even with stochasticity, the fundamental trade-off leading to an optimal CoT length persists.
Appendix G Proof
@@ G.1 Proof of Proposition 1
The second and third conclusions can be easily derived through monotonic composition, so we primarily focus on proving the first point. We begin the proof by incorporating the notation from G.2. We have
@@
Let $F(x^{*}(T),T)=g^{\prime}(x^{*}(T))=0$ We want to see how $x^{*}(T)$ changes as $T$ changes, therefore we take total derivative w.r.t. $T$. By the chain rule,
@@
According to Lemma G.1, $F(x_{0}(T),T)<0$. Since $F(x^{*}(T),T)=0,\text{ and }\,F_{x}\,\Bigl{(}x^{*}(T),\,T\Bigr{)}<0,$ we have $x_{0}(T)>x^{*}(T)$.
@@
G.5 Proof of Theorem F.2
[…]
Proof.
According to the multidimensional version of Hölder’s inequality,
$\mathbb{E}\left[\prod_{i=1}^{N}(1-e_{i})(1-\sigma_{i})\right]$ $\leq\prod_{i=1}^{N}\left(\mathbb{E}[(1-e_{i})^{2N}]\mathbb{E}[(1-\sigma_{i})^{2N}]\right)^{\frac{1}{2N}}$ (21)
(Lemma G.2) $\leq\prod_{i=1}^{N}\left(1-\frac{T}{C+2N-1}\right)\left(1-\frac{T}{NM+2N-1}\right)$ (22)
$=\left[\left(1-\frac{T}{C+2N-1}\right)\left(1-\frac{T}{NM+2N-1}\right)\right]^{N}$ (23)
G.6 Proof of Corollary 4.2
[…]
Proof.
We treat the choice of CoT length as a $k$-armed stochastic bandit with action set $\mathcal{A}=\{N_{1},\dots,N_{k}\}$ and unknown success probabilities [^3] By Proposition 1, $A(N_{i})$ is the probability that the final answer is correct when a chain of length $N_{i}$ is used. The bandit is stationary because $A(N_{i})$ does not depend on time or the agent’s past actions. $A(N_{i})\in(0,1)$. Without loss of generality, relabel the arms so that
$A(N_{1})=\max_{j}A(N_{j})\eqqcolon A^{*},\qquad A(N_{1})\geq A(N_{2})\geq\cdots\geq A(N_{k}).$
The agent uses a softmax (Gibbs) policy
$\pi_{\theta}(N_{i})\;=\;\frac{e^{\theta_{i}}}{\sum_{j=1}^{k}e^{\theta_{j}}},\qquad\theta\in\mathbb{R}^{k},$ (24)
and maximises the expected reward
$J(\theta)\;=\;\sum_{i=1}^{k}\pi_{\theta}(N_{i})\,A(N_{i}).$ (25)
Because $\pi_{\theta}$ is $C^{\infty}$ in $\theta$ and $A(N_{i})$ are constants, $J$ is smooth.
Under the REINFORCE estimator with sufficiently small, fixed step size $\eta>0$, gradient ascent updates take the form
$\theta^{(t+1)}\;=\;\theta^{(t)}+\eta\,\nabla_{\theta}J\bigl{(}\theta^{(t)}\bigr{)},$ (26)
where
$\frac{\partial J}{\partial\theta_{i}}\;=\;\pi_{\theta}(N_{i})\Bigl{(}A(N_{i})-J(\theta)\Bigr{)}.$ (27)
Eq. (27) is the classical replicator (or logit) gradient. Define the simplex $\Delta^{k-1}\!:=\!\{\pi\in(0,1]^{k}\mid\sum_{i}\pi_{i}=1\}$ and write $\pi_{\theta}=(\pi_{\theta}(N_{1}),\dots,\pi_{\theta}(N_{k}))$.
Letting $\eta\to 0$ yields the ODE
$\dot{\pi}_{i}\;=\;\pi_{i}\bigl{(}A(N_{i})-\langle\pi,A\rangle\bigr{)},\qquad i=1,\dots,k,$ (28)
with $\langle\pi,A\rangle=\sum_{j}\pi_{j}A(N_{j})$. Eq. (28) is the replicator dynamics for a fitness landscape $A$ on $\Delta^{k-1}$.
Consider the Kullback–Leibler divergence to the optimal pure strategy $\mathbf{e}_{1}=(1,0,\dots,0)$,
$V(\pi)\;=\;\sum_{i=1}^{k}\pi_{i}\ln\!\Bigl{(}\tfrac{\pi_{i}}{e_{1,i}}\Bigr{)}\;=\;-\ln\pi_{1}.$
$V$ is non-negative on $\Delta^{k-1}$ and $V(\pi)=0$ iff $\pi=\mathbf{e}_{1}$.
Taking the time derivative along Eq. (28) gives
[…]
with equality iff $\pi_{1}=1$ or $A(N_{1})=\langle\pi,A\rangle$. The latter can only happen if $\pi_{1}=1$ because $A(N_{1})>A(N_{j})$ for $j>1$. Hence $V$ is a strict Lyapunov function, and $\mathbf{e}_{1}$ is the unique asymptotically stable equilibrium of Eq. (28). All other stationary points (mixtures over sub-optimal arms) are unstable.
For sufficiently small but fixed $\eta$ (choose $\eta<\frac{1}{A^{*}}$, which always exists), projected gradient ascent is a perturbed discretisation of Eq. (28). Standard results for primal-space mirror descent imply that the discrete iterates $\pi^{(t)}\equiv\pi_{\theta^{(t)}}$ converge almost surely to the set of asymptotically stable equilibria of the ODE, i.e. to $\{\mathbf{e}_{1}\}$. Therefore
$\lim_{t\to\infty}\pi_{\theta^{(t)}}(N_{i})=\begin{cases}1,&\text{if }i=\arg\max_{j}A(N_{j}),\\[4.0pt] 0,&\text{otherwise.}\end{cases}$
Because $A$ may attain its maximum at several arms, the limit is a deterministic policy that places all probability on some maximiser of $A$.
Thus gradient ascent on Eq. (25) converges to a deterministic policy that always selects an optimal CoT length $N^{*}=\arg\max_{N\in\mathcal{A}}A(N)$, completing the proof. ∎
G.7 Technical Lemmas
Lemma G.1 (test point).
@@ G.7 Technical Lemmas
Lemma G.2 (Estimation of the $n$-th Moment of the Beta Distribution).
Let $x\sim\mathrm{Beta}(\alpha,\beta)$. Then
$\mathbb{E}[(1-x)^{n}]\leq\left(1-\frac{\alpha}{\alpha+\beta+n-1}\right)^{n}.$
Proof.
[…]
$=\frac{B(\alpha,\beta+n)}{B(\alpha,\beta)}$
$=\frac{\Gamma(\alpha)\Gamma(\beta+n)}{\Gamma(\alpha+\beta+n)}\cdot\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}$
$=\frac{\Gamma(\beta+n)}{\Gamma(\beta)}\cdot\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha+\beta+n)}$
$=\prod_{i=0}^{n-1}\frac{\beta+i}{\alpha+\beta+i}$
$\leq\left(\frac{\beta+n-1}{\alpha+\beta+n-1}\right)^{n}$
$=\left(1-\frac{\alpha}{\alpha+\beta+n-1}\right)^{n}.$
Appendix H Pseudo-code of Length-filtered Vote
Algorithm 1 Length-filtered Vote
2:Output: Final Answer $\hat{a}$
4:Define $\mathcal{A}(c)$ as the corresponding answer of candidates $c$.
7:     for $a\in\mathcal{A}$ do
$p_{j}[a]=\frac{\sum_{c\in L_{j}}\mathbb{I}(\mathcal{A}(c)=a)}{|L_{j}|}$
8:     end for
9:end for
12:return $\hat{a}$